# PO-Flow: Flow-based Generative Models for Sampling Potential Outcomes and Counterfactuals

### v1

@@ 5.4 Counterfactual Outcome Estimation
Evaluation Metrics: We evaluate the predicted counterfactual outcomes using the root mean squared error (RMSE) between the estimated counterfactuals $\hat{y}_{i,\text{cf}}^{(1-a)}$ and the ground-truth counterfactuals $y_{i,\text{cf}}^{(1-a)}$, computed as $\text{RMSE}=\small\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i,\text{cf}}^{(1-a)}-\hat{y}_{i,\text{cf}}^{(1-a)})^{2}}$. The results are presented in Table 5. Here, $\text{RMSE}_{\text{in}}$ and $\text{RMSE}_{\text{out}}$ denote the RMSE evaluated on the training and test sets, respectively.
Baselines: To our knowledge, very few existing methods are specifically designed for counterfactual outcome prediction conditioned on factuals in general settings—without relying on structural causal models or temporal information. While methods like CFR include “counterfactual” in name, they target unobserved potential outcomes rather than factual-conditioned ones, though we still include them as valuable baselines. We also adapt DiffPO and GANITE for this task, as detailed in Appendix B.6.
@@
[…]
@@
5.5 Counterfactual Image Generation
Figure 3: PO-Flow counterfactual image generation from female to male faces. $(a)$: Each row shows a factual outcome $y^{(a)}$ (1st column), its encoded latent $z_{y}$ (2nd column), and reverse-time decoding at $t=2/3$, $t=1/3$, and $t=0$ (3rd–5th columns), yielding the counterfactual $\hat{y}_{\text{cf}}^{(1-a)}$. $(b)$: Counterfactuals decoded from perturbed latents $z_{y}+\delta$, with $\delta\sim\mathcal{N}(0,0.2I)$.
Beyond low-dimensional datasets, PO-Flow also succeeds on high-dimensional counterfactual image generation. We train PO-Flow on the latent space of a pre-trained VAE using the CelebA dataset, predicting counterfactual male faces from original female faces. In this setting, $X\in\mathbb{R}^{6}$ encodes visual attributes (hair, eyebrows, bangs, lips), $a\in\{\text{male},\text{female}\}$ denotes gender, and $y\in\mathbb{R}^{512}$ is the VAE latent feature. Visual results are shown in Figures 3 and 6; detailed settings are in Appendix B.2.
@@ 5.6 Discussions and Comparisons
Training efficiency: PO-Flow exhibits high training efficiency, converging in approximately 300 gradient steps across multiple causal datasets, as shown in Appendix C.1.
Minimal architecture sensitivity: We find that PO-Flow is the least sensitive to neural network architecture and hyperparameters. Using the same NN structure across all causal datasets, PO-Flow consistently achieves state-of-the-art performance. In contrast, we find that generative approaches including DiffPO, INFs, and GANITE require careful hyperparameter tuning for each dataset.
Unique density estimation: Our continuous normalizing flow (CNF) backbone enables density learning of potential outcomes, as shown in Proposition 4.2 and Section 5.3.1, which, to our knowledge, is the first generative model in the field to achieve this.
Counterfactual: We take a step further by exploring counterfactual outcomes, enabled by the smooth PO-Flow mapping that preserves factual information during encoding. Unlike diffusion-based DiffPO, which has noisy forward processes, PO-Flow offers information-preserving forward processes, as shown in Figure 1. To our knowledge, it is among the first to explore empirical counterfactual prediction without relying on temporal data or explicitly modeling structural causal models.
However, one area for improving counterfactual prediction lies in modeling uncertainty more expressively. The current covariate-conditioned Gaussian perturbations may not fully capture counterfactual variation. Future work could explore more theoretically grounded counterfactual uncertainty.
@@ 6 Conclusions
We introduced PO-Flow, a continuous normalizing flow framework for modeling potential and counterfactual outcomes in causal inference. PO-Flow provides a unified generative framework for potential outcome prediction, CATE estimation, and general-purpose counterfactual prediction—without explicitly modeling structural causal models (SCMs) or relying on auxiliary information for the counterfactual task.
Empirically, PO-Flow achieves state-of-the-art performance on benchmark datasets across multiple causal tasks. Among generative models, it is the first to accurately learn the density of potential outcomes, enabling principled uncertainty quantification. Furthermore, PO-Flow is efficient, robust, and extendable to high-dimensional settings such as counterfactual image generation. These results highlight PO-Flow as a unified and scalable solution for causal inference tasks.
@@ Appendix A Proofs
Proposition 4.1. Assume that $p_{t}(y|x,a)>0$ for all $y\in\mathbb{R}^{d_{Y}}$ and $t\in[0,1]$. Then, up to a constant independent of $\theta$, the Conditional Flow Matching (CFM) loss and the original Flow Matching (FM) loss are equivalent. Hence,
@@
We denote the conditional reference velocity field as $u(y,t,x,a|y_{0})=\frac{d\phi_{t}}{dt}=(1-\sigma_{\text{min}})y_{1}-y_{0}$, as defined in (6). We further define the marginal (unconditional) reference velocity field as:
[…]
@@
$\mathcal{L}_{\text{FM}}=\mathbb{E}_{t\sim\mathcal{U}[0,1],p_{Y_{t},X,A}(\cdot)}\|v_{\theta}(y_{t},t,x,a)-u(y_{t},t,x,a)\|^{2}.$ (15)
In the main text, we chose $\phi_{t}$ as the linear interpolant between $y_{0}$ and $y_{1}$, which results in a constant reference velocity $\frac{d\phi_{t}}{dt}$ independent of $t$. This leads to the simplified CFM loss presented in Eq. (9). However, the general CFM loss, corresponding to an arbitrary reference velocity field $u(y,t\mid y_{0})$ conditioned on $y_{0}$, is defined as:
$\mathcal{L}_{\text{CFM}}=\mathbb{E}_{t\sim\mathcal{U}[0,1],p_{Y}(y_{0}),p_{Y_{t},X,A|Y_{0}}(\cdot|y_{0})}\|v_{\theta}(\cdot)-u_{t}(\cdot|y_{0})\|^{2}$ (16)
@@
Besides, the term $\mathbb{E}_{t\sim\mathcal{U}[0,1],p_{Y_{t},X,A}(\cdot)}\|v_{\theta}(\cdot)\|^{2}$ can be equivalently expressed in conditional form as:
[…]
$=\int\|v_{\theta}(\cdot)\|^{2}p_{Y_{t},X,A|Y_{0}}(\cdot|y_{0})p_{Y}(y_{0})\,dy_{0}\,dy\,dx\,da\,dt$
$=\mathbb{E}_{t\sim\mathcal{U}[0,1],p_{Y}(y_{0}),p_{Y_{t},X,A|Y_{0}}(\cdot|y_{0})}\|v_{\theta}(\cdot)\|^{2}.$
@@
[…]
$=\int\left\langle v_{\theta}(\cdot),\int u(\cdot|y_{0})p_{Y_{t},X,A}(\cdot|y_{0})p_{Y}(y_{0})\,dy_{0}\right\rangle\,dy\,dx\,da\,dt$
$=\int\langle v_{\theta}(\cdot),u(\cdot|y_{0})\rangle p_{Y_{t},X,A}(\cdot|y_{0})p_{Y}(y_{0})\,dy_{0}\,dy\,dx\,da\,dt$
$=\mathbb{E}_{t\sim\mathcal{U}[0,1],p_{Y}(y_{0}),p_{Y_{t},X,A|Y_{0}}(\cdot|y_{0})}\langle v_{\theta}(\cdot),u(\cdot|y_{0})\rangle.$
Finally, because $u_{t}(\cdot\mid y_{0})$ does not depend on $\theta$, substituting Eqs. (18) and (19) into Eq. (15) and differentiating with respect to $\theta$ yields:
@@
Proposition 4.2. Given a noise sample $y_{1}\sim q(\cdot)$ at $t=1$, the log-density of the resulting potential outcome at $t=0$, obtained via PO-Flow, is given by:
$\log\hat{p}_{Y}(\cdot|x,a,y_{1})=\log q(y_{1})+\int_{0}^{1}\nabla\cdot v_{\theta}(y(t),t,x,a)dt.$
@@
$\frac{dy(t)}{dt}=v_{\theta}(y(t),t,x,a),\quad t\in[0,1],$ (21)
@@
$\partial_{t}p(y,t)+\nabla\cdot\big{(}p(y,t)v_{\theta}(y,t,x,a)\big{)}=0.$ (22)
Here, $p(y,t)$ denotes the time-dependent probability density of the samples $y$.
Next, we have that the dynamics of the density $p(\cdot)$ governed by the velocity field $v_{\theta}(y,t,x,a)$ is given by:
$\frac{d}{dt}\log p(y(t),t)$ $=\frac{\nabla p(y(t),t)\cdot\partial_{t}y(t)+\partial_{t}p(y(t),t)}{p(y(t),t)}$ (23)
$=\frac{\nabla p\cdot\partial_{t}y+\partial_{t}p}{p}\Big{|}_{(y(t),t)}$ (24)
@@
$=\frac{\nabla p\cdot v_{\theta}-(\nabla p\cdot v_{\theta}+p\nabla\cdot v_{\theta})}{p}\Big{|}_{(y(t),t)}$ (26)
@@
Starting from an initial sample $y_{1}\sim q(\cdot)$ and integrating from $t=1$ to $t=0$, we have:
$\log\hat{p}_{Y}(\cdot|x,a,y_{1})=\log q(y_{1})+\int_{0}^{1}\nabla\cdot v_{\theta}(y(t),t,x,a)dt.$ (28)
@@
$\log\hat{p}_{Y}(\cdot|x,a,y_{1})=-\frac{d_{Y}}{2}\log 2\pi-\frac{1}{2}\|y_{1}\|^{2}+\int_{0}^{1}\nabla\cdot v_{\theta}(y(t),t,x,a)dt.$ (29)

### v2

@@ 5.4 Counterfactual Outcome Estimation
Evaluation Metrics: We assess counterfactual predictions using the RMSE between model-estimated $\hat{y}_{\text{cf}}^{(1-a)}$ and ground-truth $y_{\text{cf}}^{(1-a)}$. Results are in Table 6.
Baselines: To our knowledge, very few existing methods are specifically designed for counterfactual outcome prediction conditioned on factuals in general settings—without explicitly modeling structural causal models or relying on temporal information. While methods like CFR include “counterfactual” in name, they target unobserved potential outcomes rather than factual-conditioned ones, though we still include them as valuable baselines. We also adapted DiffPO and GANITE for this task, as detailed in Appendix B.5.
@@
PO-Flow $\mathbf{1.50_{\pm.15}}$ $\mathbf{1.53_{\pm.17}}$ $\mathbf{0.04_{\pm.01}}$ $\mathbf{0.05_{\pm.02}}$ $\mathbf{1.63_{\pm.09}}$ $1.69_{\pm.10}$ $\mathbf{0.04_{\pm.01}}$ $\mathbf{0.04_{\pm.02}}$
@@ 5.5 Discussions and Comparisons
Training efficiency: PO-Flow exhibits high training efficiency, converging in approximately 300 gradient steps across multiple causal datasets (Appendix D.3), and demonstrates notable advantages in both training and sampling time compared to other baselines (Table 9).
Density estimation for principled uncertainty quantification: Our continuous normalizing flow (CNF) backbone enables density learning of potential outcomes, as detailed in Section 5.3.1. This property allows us to generate multiple potential outcomes for each individual and to identify the most likely ones based on their log-densities. The empirical evaluation of this procedure is presented in Table 4.
Counterfactual: We take a step further by exploring counterfactual outcomes, enabled by the smooth PO-Flow mapping that preserves factual information during encoding. Under (A1)–(A3), we include a supporting recovery result for this architecture (Proposition 12; Corollary 4.5). It is among the first to explore counterfactual outcomes in the potential outcomes framework without relying on temporal data or explicitly modeling structural causal models.
@@ 6 Conclusions
We introduced PO-Flow, a continuous normalizing flow framework for modeling potential and counterfactual outcomes in causal inference. PO-Flow provides a unified generative framework for potential outcome prediction, CATE estimation, and general-purpose counterfactual prediction in potential outcomes—without explicitly modeling structural causal models (SCMs) or relying on auxiliary information. To our knowledge, this constitutes one of the earliest attempts to address counterfactual outcome prediction in the potential outcomes literature, a crucial yet underexplored direction.
Theoretically, we include a supporting recovery result for the proposed architecture (Corollary 4.5). Empirically, PO-Flow attains state-of-the-art performance on benchmark datasets across standard causal tasks, while additionally supporting counterfactual outcome prediction and log-density estimation for principled uncertainty quantification. In addition, PO-Flow is both computationally efficient and robust compared to modern baselines. These results highlight PO-Flow as a unified and scalable solution for potential-outcomes–based causal tasks, supporting reliable decision-making in domains such as healthcare.
@@ Appendix A Proofs
Proposition A.1.
Assume that $p(y|x,a)>0$ for all $y$ and $t\in[0,1]$. Then, up to a constant independent of $\theta$, the Conditional Flow Matching (CFM) loss and the original Flow Matching (FM) loss are equivalent. Hence,
@@
We denote the conditional reference velocity field as $u(y,t,x,a|y_{0})=\frac{d\phi}{dt}=(1-\sigma_{\text{min}})y_{1}-y_{0}$, as defined in (6). We further define the marginal (unconditional) reference velocity field as:
$u(y,t,x,a)=\int u(y,t,x,a|y_{0})\frac{p(y,t,x,a|y_{0})p_{Y}(y_{0})}{p(y,t,x,a)}dy_{0},$ (15)
@@
$\mathcal{L}_{\text{FM}}=\mathbb{E}_{t\sim\mathcal{U}[0,1],p_{Y,X,A}(\cdot)}\|v_{\theta}(y,t,x,a)-u(y,t,x,a)\|^{2}.$ (16)
In the main text, we chose $\phi$ as the linear interpolant between $y_{0}$ and $y_{1}$, which results in a constant reference velocity $\frac{d\phi}{dt}$ independent of $t$. This leads to the simplified CFM loss presented in Eq. (9). However, the general CFM loss, corresponding to an arbitrary reference velocity field $u(y,t\mid y_{0})$ conditioned on $y_{0}$, is defined as:
$\mathcal{L}_{\text{CFM}}=\mathbb{E}_{t\sim\mathcal{U}[0,1],p_{Y}(y_{0}),p_{Y,X,A|Y_{0}}(\cdot|y_{0})}\|v_{\theta}(\cdot)-u(\cdot|y_{0})\|^{2}$ (17)
@@
Besides, the term $\mathbb{E}_{t\sim\mathcal{U}[0,1],p_{Y,X,A}(\cdot)}\|v_{\theta}(\cdot)\|^{2}$ can be equivalently expressed in conditional form as:
$\mathbb{E}_{t\sim\mathcal{U}[0,1],p_{Y,X,A}(\cdot)}\|v_{\theta}(\cdot)\|^{2}$ $=\int\|v_{\theta}(\cdot)\|^{2}p_{Y,X,A}(\cdot)\,dy\,dx\,da\,dt$ (19)
$=\int\|v_{\theta}(\cdot)\|^{2}p_{Y,X,A|Y_{0}}(\cdot|y_{0})p_{Y}(y_{0})\,dy_{0}\,dy\,dx\,da\,dt$
$=\mathbb{E}_{t\sim\mathcal{U}[0,1],p_{Y}(y_{0}),p_{Y,X,A|Y_{0}}(\cdot|y_{0})}\|v_{\theta}(\cdot)\|^{2}.$
@@
$\mathbb{E}_{t\sim\mathcal{U}[0,1],p_{Y,X,A}(\cdot)}\langle v_{\theta}(\cdot),u(\cdot)\rangle$ $=\int\left\langle v_{\theta}(\cdot),\frac{\int u(\cdot|y_{0})p_{Y,X,A}(\cdot|y_{0})p_{Y}(y_{0})\,dy_{0}}{p_{Y,X,A}(\cdot)}\right\rangle p_{Y,X,A}(\cdot)\,dy\,dx\,da\,dt$ (20)
$=\int\left\langle v_{\theta}(\cdot),\int u(\cdot|y_{0})p_{Y,X,A}(\cdot|y_{0})p_{Y}(y_{0})\,dy_{0}\right\rangle\,dy\,dx\,da\,dt$
$=\int\langle v_{\theta}(\cdot),u(\cdot|y_{0})\rangle p_{Y,X,A}(\cdot|y_{0})p_{Y}(y_{0})\,dy_{0}\,dy\,dx\,da\,dt$
$=\mathbb{E}_{t\sim\mathcal{U}[0,1],p_{Y}(y_{0}),p_{Y,X,A|Y_{0}}(\cdot|y_{0})}\langle v_{\theta}(\cdot),u(\cdot|y_{0})\rangle.$
Finally, because $u(\cdot\mid y_{0})$ does not depend on $\theta$, substituting Eqs. (19) and (20) into Eq. (16) and differentiating with respect to $\theta$ yields:
@@
Proposition 12. (Encoded $Z$ as a function of the exogenous noise $U$) Let Assumption 4.1 hold. Assume that the exogenous noise $U\sim\mathrm{Unif}[0,1]$. Then there exists a continuously differentiable bijection function $\psi_{X}:\;\mathcal{U}\;\longrightarrow\;\mathcal{Z}$ that is independent of the treatment assignment $A$, i.e.,
$Z\;=\;\Phi_{\theta}(Y^{(A)};\ X,A)\;=\;\Phi_{\theta}(f_{X}(A,U);X,A)\;=\;\psi_{X}\bigl(U\bigr)\quad\text{a.s.}$ (22)
Proof. Based on the structural causal model, we have $Y^{(A)}=f_{X}(A,U)$. We define
$m_{X,A}(U)\;:=Z=\Phi_{\theta}(Y^{(A)};\,X,A)=\;\Phi_{\theta}\bigl(f_{X}(A,U);\ X,A\bigr),$
as a function that may depend on the action $A$. Therefore, our target becomes proving that $m_{X,A}(U)$ is functionally invariant to $A$.
By (A3) in Assumption 4.1, we have $Z\perp\!\!\!\perp A\ |\ X$. Therefore,
$p_{\theta,Z\mid X,A}(z)=p_{\theta,Z}(z).$ (23)
In continuous normalizing flows, $\Phi_{\theta}$ is invertible with respect to $Y^{(A)}$ due to the invertibility of the underlying neural ODE. In 1-D, this implies that $\Phi_{\theta}$ is monotonic in $Y^{(A)}=f_{X}(A,U)$. Without loss of generality, we assume that $\Phi_{\theta}$ is strictly increasing in $f_{X}$.
Moreover, by (A2), since $f_{X}(\cdot,U)$ is strictly monotone in $U$. Without loss of generality, we assume strictly increasing. Then it follows by the composition rule that $m_{X,A}(U)$ is strictly increasing in $U$ and hence bijective on $[0,1]$.
Since $Z=\Phi_{\theta}(Y^{(A)};\,X,A)=m_{X,A}(U)$, we may apply change of variables formula w.r.t. $U$:
$p_{\theta,Z|X,A}(z)\;=\;p_{U}\!\bigl(m_{X,A}^{-1}(z)\bigr)\;\left|\frac{d}{dz}m_{X,A}^{-1}\!\bigl(z\bigr)\right|=1\cdot\frac{d}{dz}m_{X,A}^{-1}\!\bigl(z\bigr),$ (24)
where the last equation follows from the uniform distribution of $U$ and the fact that $p_{Z}>0$.
Then based on (23), we have:
$p_{\theta,Z|X,A=1}(z)=\frac{d}{dz}m_{X,1}^{-1}\!\bigl(z\bigr)=\frac{d}{dz}m_{X,0}^{-1}\!\bigl(z\bigr)=p_{\theta,Z|X,A=0}(z).$ (25)
It follows that:
$m_{X,A}^{-1}(z)=\int^{z}\frac{d}{dt}m_{X,A}^{-1}\!\bigl(t\bigr)dt=\int^{z}c(t)dt+c_{A},$ (26)
where $c(t)=\frac{d}{dt}m_{X,1}^{-1}\!\bigl(t\bigr)=\frac{d}{dt}m_{X,0}^{-1}\!\bigl(t\bigr)$ is independent of $A$, and $c_{A}$ is a constant for each $A$.
By re-inverting (26), we have:
$m_{X,A}(u)=(m_{X,A}^{-1})^{-1}(u)=\Bigl(\underbrace{\int^{\bullet}c(t)\,\mathrm{d}t}_{:=H(\bullet)}+c_{A}\Bigr)^{-1}(u)=H^{-1}\bigl(u-c_{A}\bigr).$ (27)
Since $m_{X,A}$ is a bijection $0,1]\to\text{supp}(Z)=\{z\in\mathbb{R}:p_{Z}(z)>0\}$, we have:
$m_{X,A}(0)=\inf\ \text{supp}(Z),\ \ m_{X,A}(1)=\sup\ \text{supp}(Z),\ \text{for both}\ A.$ (28)
Besides, since the support of $Z$ does not depend on $A$ because of (23), we have that
$\inf\ \text{supp}(Z)=m_{X,0}(0)=H^{-1}(-c_{0})=H^{-1}(-c_{1})=m_{X,1}(0).$ (29)
Therefore, $c_{0}=c_{1}:=c$ is a constant that does not depend on $A$.
Therefore, we can write
$\psi_{X}(U)\;:=\;H^{-1}\bigl(U-c\bigr)\;=\;m_{X,0}(U)\;=\;m_{X,1}(U).$ (30)
As a result, we conclude that
$Z\;=\;\Phi_{\theta}\bigl(Y^{(A)};\,X,A\bigr)\;=\;\psi_{X}\bigl(U\bigr)\quad\text{a.s.},$ (31)
is a function that only depends on the exogenous noise $U$ (and covariates $X$), and is invariant to the treatment $A$.
Corollary 4.5 (Counterfactual recovery) Assume the conditions in Assumption 4.1 hold. Consider a factual sample $\bigl(Y^{(A)},X,A\bigr)$ given by the structural equation $Y^{(A)}=f_{X}\bigl(A,U\bigr)$, and let its encoded representation be $Z=\Phi_{\theta}(Y^{(A)};X,A)$. Apply the intervention $\mathrm{do}\bigl(A=\gamma\bigr)$. Assume that the true counterfactual is given by $Y_{\mathrm{CF}}^{(\gamma)}=f_{X}(\gamma,U)$.
The decoder then recovers the counterfactual almost surely:
$\hat{Y}_{\mathrm{CF}}^{\gamma}:=\Phi_{\theta}^{-1}\bigl(Z;\ X,\gamma\bigr)=Y_{\mathrm{CF}}^{\gamma}.$ (32)
Proof. Due to the deterministic and invertible nature of the continuous normalizing flow when the conditioning input is fixed, we have,
$\Phi_{\theta}^{-1}(\Phi_{\theta}(Y;\,X,A);X,A)=Y.$ (33)
From Proposition 4.3, we have $Z=\;\Phi_{\theta}\bigl(Y^{(A)};\,X,A\bigr)\;=\Phi_{\theta}(f_{X}(A,U);\ X,A\bigr)\;=\;\psi_{X}\bigl(U\bigr)$ does not depend on $A$. Therefore, fixing $U$ as the same and plugging in $\mathrm{do}\bigl(A=\gamma\bigr)$, we have:
$\Phi_{\theta}\bigl(Y^{(A)};\,X,A\bigr)=\Phi_{\theta}\bigr(f_{X}(A,U);\ X,A\bigr)=\psi_{X}\bigl(U\bigr)=\Phi_{\theta}\bigr(f_{X}(\gamma,U);\ X,\gamma\bigr)=\Phi_{\theta}(Y_{\mathrm{CF}}^{\gamma};\ X,\gamma),$ (34)
or more simply,
$\Phi_{\theta}\bigl(Y^{(A)};\,X,A\bigr)=\Phi_{\theta}(Y_{\mathrm{CF}}^{\gamma};\ X,\gamma),$ (35)
we further have,
$\Phi_{\theta}^{-1}(\Phi_{\theta}\bigl(Y^{(A)};\,X,A\bigr);\ X,\gamma)=\Phi_{\theta}^{-1}(\Phi_{\theta}(Y_{\mathrm{CF}}^{\gamma};\ X,\gamma);\ X,\gamma).$ (36)
Noting that the L.H.S. of (36) is exactly the counterfactual procedures (encoding-decoding) of PO-Flow, and by (33), the R.H.S. of (36) equals to $Y_{\mathrm{CF}}^{\gamma}$. Therefore,
$\hat{Y}_{\mathrm{CF}}^{\gamma}:=\Phi_{\theta}^{-1}\bigl(Z;\ X,\gamma\bigr)=Y_{\mathrm{CF}}^{\gamma}.$ (37)
Proposition 4.7. Given a noise sample $z\sim q(\cdot)$ at $t=1$, the log-density of the resulting potential outcome at $t=0$, obtained via PO-Flow, is given by:
$\log p_{\theta,Y\mid X,A}\bigl(y(0)\mid x,a,z\bigr)=\log q(z)\;+\;\int_{0}^{1}\nabla_{y}\!\cdot\,v_{\theta}\bigl(y(t),t;\,x,a\bigr)\,dt.$
@@
$\frac{dy(t)}{dt}=v_{\theta}(y(t),t;x,a),\quad t\in[0,1],$ (38)
@@
$\partial p_{\theta}(y,t)+\nabla\cdot\big(p_{\theta}(y,t)v_{\theta}(y,t;x,a)\big)=0.$ (39)
Here, $p_{\theta}(y,t)$ denotes the time-dependent probability density of the samples $y$.
Next, we have that the dynamics of the density $p_{\theta}(\cdot)$ governed by the velocity field $v_{\theta}(y,t,x,a)$ is given by:
$\frac{d}{dt}\log p_{\theta}(y(t),t)$ $=\frac{\nabla p_{\theta}(y(t),t)\cdot\partial_{t}y(t)+\partial_{t}p_{\theta}(y(t),t)}{p_{\theta}(y(t),t)}$ (40)
$=\frac{\nabla p_{\theta}\cdot\partial_{t}y+\partial_{t}p_{\theta}}{p_{\theta}}\Big|_{(y(t),t)}$ (41)
@@
$=\frac{\nabla p_{\theta}\cdot v_{\theta}-(\nabla p_{\theta}\cdot v_{\theta}+p_{\theta}\nabla\cdot v_{\theta})}{p_{\theta}}\Big|_{(y(t),t)}$ (43)
@@
Starting from an initial sample $z\sim q(\cdot)$ and integrating from $t=1$ to $t=0$, we have:
$\log p_{\theta,Y\mid X,A}\bigl(y(0)\mid x,a,z\bigr)=\log q(z)\;+\;\int_{0}^{1}\nabla_{y}\!\cdot\,v_{\theta}\bigl(y(t),t;\,x,a\bigr)\,dt.$ (45)
@@
$\log p_{\theta,Y\mid X,A}\bigl(y(0)\mid x,a,z\bigr)=-\frac{d_{Y}}{2}\log 2\pi-\frac{1}{2}\|z\|^{2}+\int_{0}^{1}\nabla_{y}\cdot v_{\theta}(y(t),t;x,a)dt.$ (46)